OpenStudy (anonymous):

What statement makes the open sentence 12 + 3x = 30 true? A. x = 2 B. x = 4 C. x = 5 D. x = 6

OpenStudy (anonymous):

D

OpenStudy (anonymous):

12 + 3*2= 12 + 3*4= 12+ 3*5= 12+ 3*6=

OpenStudy (anonymous):

add and multiply those and you will get your answer